A recently hired chief executive officer wants to reduce future production costs to improve the company’s earnings, thereby increasing the value of the company’s stock. The plan is to invest $96,000 now and $60,000 in each of the next 4 years to improve productivity. By how much must annual costs decrease in years 5 through 15 to recover the investment plus a return of 13% per year? The annual cost decreases by $
Answer #1

Answer:

Let x be the annual decreases in cost.

r=15%

Initial Investment P=$96,0000

Annual investment A=$40,000

to recover investment

PV of investment =PV of Annual Decrease in cost

P+A*(1-(1+r)^-4)/r=x*(1-(1+r)^-11)/{r*(1+r)^4}

96000+40000*(1-(1+15%)^-4)/15%=X*(1-(1+15%)^-11)/{15%*(1+15%)^4}

X=$70244.52